# Beyond Chemical QA: Evaluating LLM’s Chemical Reasoning with Modular Chemical Operations

### v2

@@ Abstract
While large language models (LLMs) with Chain-of-Thought (CoT) reasoning excel in mathematics and coding, their potential for systematic reasoning in chemistry, a domain demanding rigorous structural analysis for real-world tasks like drug design and reaction engineering, remains untapped. Current benchmarks focus on simple knowledge retrieval, neglecting step-by-step reasoning required for complex tasks such as molecular optimization and reaction prediction. To address this, we introduce ChemCoTBench, a reasoning framework that bridges molecular structure understanding with arithmetic-inspired operations, including addition, deletion, and substitution, to formalize chemical problem-solving into transparent, step-by-step workflows. By treating molecular transformations as modular "chemical operations", the framework enables slow-thinking reasoning, mirroring the logic of mathematical proofs while grounding solutions in real-world chemical constraints. We evaluate models on two high-impact tasks: Molecular Property Optimization and Chemical Reaction Prediction. These tasks mirror real-world challenges while providing structured evaluability. By providing annotated datasets, a reasoning taxonomy, and baseline evaluations, ChemCoTBench bridges the gap between abstract reasoning methods and practical chemical discovery, establishing a foundation for advancing LLMs as tools for AI-driven scientific innovation. 111 * Equal Contribution; $\dagger$ Corresponding Authors. 222 ChemCoTBench: https://huggingface.co/datasets/OpenMol/ChemCoTBench 333 ChemCoTDataset: https://huggingface.co/datasets/OpenMol/ChemCoTDataset
@@ 1 Introduction
To address these limitations, we introduce ChemCoTBench, a step-by-step, application-oriented, and high-quality benchmark for evaluating LLM reasoning in chemical applications. A core innovation of ChemCoTBench is its formulation of complex chemical tasks, specifically targeting molecular modeling and design (Fig.1), into explicit sequences of verifiable modular chemical operations on SMILES structures (e.g., substructure addition, deletion, or substitution). This approach allows for a granular assessment of an LLM’s ability to execute and chain together fundamental chemical transformations. The benchmark features progressively challenging tasks, spanning from basic molecular understanding and editing to property-guided structure optimization and complex multi-molecule chemical reactions. High-quality evaluation is ensured through a dual validation process combining LLM judgment with expert review from 13 chemists.
We employ quantitative assessments for all subtasks in ChemCoTBench to evaluate the chemical reasoning ability across reasoning-enhanced and non-reasoning LLMs. Experimental results reveal room for improvement in reasoning LLMs, particularly open-source and distilled-reasoning LLMs, when addressing complex chemical problems. While these models demonstrate strong performance in complex mathematical and coding tasks, they are unable to organize chemical knowledge and establish step-wise modular chemical operations due to the scarcity of chemical reasoning data. Notably, ChemCoTDataset, the large chemical CoT dataset provided by ChemCoTBench, is shown to enhance chemical reasoning performance, effectively addressing the reasoning data scarcity issue in Chemistry.
@@ 2 Related Works
LLM Chain-of-Thoughts.  LLMs have progressed from text generators to reasoning systems, with [64]’s Chain-of-Thought enabling stepwise problem decomposition via "slow-thinking" paradigms. These reasoning-enhanced LLMs have shown impressive performance in domains requiring systematic problem-solving skills, particularly in mathematics and coding. Models like DeepSeek-R1 [13], Gemini [56], and Anthropic Claude [53] have achieved notable results on mathematical benchmarks like MATH [19] and GSM8K [6], while also excelling at programming. Recent studies have begun exploring LLMs for chemical tasks, such as synthesis planning [4] and computational chemistry [45, 51, 26]. However, these efforts lack a systematic evaluation of LLMs’ chemical reasoning capabilities, spanning spatial reasoning, domain-specific knowledge integration, and multi-step logical inference.
@@ 3 ChemCoTBench Construction
ChemCoTBench contains 1,495 samples across 22 chemical tasks as the benchmark dataset, as shown in Fig 2(a). 14,000 high-quality samples with chain-of-thoughts annotations are further sampled to form the ChemCoTDataset. ChemCoTBench was constructed through over 1,800 hours of combined expert and LLM-assisted annotation. It comprises four main tasks and 22 subtasks, covering a broad spectrum of chemical challenges. We define the reasoning steps of each task as modular chemical operations, as shown in the bottom two lines of Fig. 3. ChemCoTBench is guided by two core principles: Diversity and Quality. Molecular diversity is ensured by systematically selecting compounds with varied scaffolds and functional groups, enabling broad coverage of real-world chemical scenarios. To ensure high data quality, all benchmark samples undergo multi-stage hybrid review by LLMs and expert chemists, with prompt templates iteratively refined to meet subtask-specific requirements.
@@ 4.2 Evaluated LLMs
Our evaluation includes three model categories: (1) Reasoning LLMs with explicit step-by-step reasoning, including Deepseek-R1 [13], o1-mini [58], o3-mini [59], Gemini-2.5-pro [55], Claude-3.7-Sonnet-thinking [53], Qwen-3-thinking [60], Llama-Nemotron-thinking [2]; (2) General-purpose non-reasoning LLMs without specialized reasoning mechanisms including GPT-4o [24], Qwen-2.5/3 [66], Llama-3.3 [12], Gemma-2 [57], Phi-4 [1], OLMo2 [44] (3) Biomolecular LLMs BioMedGPT [39], BioMistral [31], and Text+Chem T5 [5]. This comprehensive comparison evaluates whether reasoning-specific capabilities provide advantages over domain-specific models in challenging chemical reasoning tasks. Details of evaluation implementation and prompt design are available in Appendix C.2.
@@
[…]
@@
[…]
Claude3.7-sonnet-think 0.41 81 0.59 77 0.09 73 0.18 66 -0.01 49 0.01 57
DeepSeek-R1 0.36 74 1.48 97 0.05 72 0.10 62 -0.06 29 -0.02 41
o3-mini@20250103 0.29 68 1.15 85 0.17 86 0.18 69 -0.08 23 -0.03 45
@@
Qwen3-235B-A22B-think -0.01 41 0.27 42 0.01 24 0.03 31 -0.01 23 0.01 31
Qwen3-32B-think 0.0 2 0.11 23 0.02 14 0.0 6 -0.02 6 -0.02 5
Llama-Nemo-49B-think -0.64 24 0.20 24 -0.16 41 -0.05 30 -0.15 7 -0.12 11
@@
[…]
DeepSeek-V3 0.08 34 0.47 93 0.08 46 0.02 28 0.0 18 0.0 29
Gemini-2.0-flash 0.35 75 0.19 54 0.10 79 0.15 63 0.03 34 0.0 38
Qwen3-235B-A22B 0.02 41 0.51 45 0.01 26 0.01 31 -0.01 23 0.0 34
Qwen3-32B -0.03 2 0.17 23 0.02 14 -0.01 6 -0.02 6 -0.02 5
Qwen2.5-72B-Instruct -0.12 42 0.28 60 0.03 57 0.04 40 -0.02 26 -0.01 40
Qwen2.5-32B-Instruct 0.03 47 0.42 66 -0.01 54 0.04 32 -0.04 19 -0.02 31
Llama-3.3-70B-Instruct -0.16 42 0.61 80 0.07 61 -0.02 31 -0.04 30 -0.02 40
Llama-Nemo-Super-49B -0.14 27 0.31 41 0.02 50 -0.02 18 -0.04 16 -0.03 27
Gemma-2-27b-it -0.03 34 0.34 66 0.05 56 -0.03 15 0.0 16 -0.01 17
Phi-4-14B -0.10 45 0.28 54 0.11 74 -0.04 18 -0.05 14 -0.04 22
OLMo2-32B-Instruct -2.03 22 1.03 46 -0.13 40 -0.11 7 -0.12 8 -0.11 12
@@ 4.3 LLMs’ Performance on Solving ChemCoTBench
4.4 Evaluating Strategies to Enhance Chemical Reasoning in Open-Source LLMs
@@
Our preceding analyses underscored the critical role of advanced reasoning capabilities (or "slow thinking") for tackling complex chemical tasks. This motivates an investigation into efficient methods for bolstering these capabilities within open-source LLMs.
@@
Challenges in Distilling Chemical Reasoning: Distilling CoT capabilities from advanced LLMs (e.g., using DeepSeek-R1-generated samples [71, 13]) is a common strategy to enhance reasoning in smaller models. However, this approach proves significantly limited for specialized chemical reasoning. Our experiments (Fig.4) show that Qwen2.5-Instruct models distilled for CoT exhibit little to no improvement on ChemCoTBench chemical subtasks compared to their non-distilled counterparts; indeed, smaller base models (1.5B-32B) often perform comparably or better. While effective for general domains like code and math (Fig.4), this distillation strategy falters in chemistry, likely due to insufficient volume or specificity of chemical CoT samples in the distillation process, hindering the development of robust step-by-step chemical reasoning. Moreover, smaller distilled models (<7B) frequently produce lengthy, repetitive, and irrelevant (hallucinatory) thought processes. These findings suggest that direct CoT distillation, without substantial domain-specific adaptation, is an ineffective standalone method for improving chemical reasoning in open-source models.
Domain-Specific Chemical CoT Data Augmentation Boosts Reasoning: Given the limitations of direct distillation, we explored enhancing chemical reasoning using our high-quality, domain-specific ChemCoTDataset via prompting. This dataset was meticulously curated to minimize hallucinations and align with expert thought processes, which we posited would be vital for chemical reasoning tasks. We tested this by evaluating two CoT prompting strategies: one providing only coarse strategic guidance (CoT templates), and another augmented with detailed step-by-step reasoning processes from our dataset. The results in the bottom line of Fig. 4 consistently demonstrate that our large-scale chemical CoT dataset significantly enhances the chemical reasoning capabilities of Qwen-2.5 models across various scales (1.5B to 32B) when used in this way. Augmentation with detailed CoT processes yielded stable and substantial performance gains across all evaluated tasks. Notably, while CoT templates stably benefited small models (<14B), larger LLMs develop rigid code & math reasoning templates that constrain their ability to effectively leverage the chemical patterns for performance improvement. This highlights an interplay between model scale and the granularity of CoT guidance required for optimal benefit.
@@
Figure 4: The top two rows compare the reasoning performance of the Qwen-2.5-Instruct series against its DeepSeek-R1-distilled versions. The bottom row illustrates performance improvements in Qwen-2.5-Instruct when enhanced with the CoT template and detailed CoT process.

### v3

@@ Abstract
While large language models (LLMs) with Chain-of-Thought (CoT) reasoning excel in mathematics and coding, their potential for systematic reasoning in chemistry, a domain demanding rigorous structural analysis for real-world tasks like drug design and reaction engineering, remains untapped. Current benchmarks focus on simple knowledge retrieval, neglecting step-by-step reasoning required for complex tasks such as molecular optimization and reaction prediction. To address this, we introduce ChemCoTBench, a reasoning framework that bridges molecular structure understanding with arithmetic-inspired operations, including addition, deletion, and substitution, to formalize chemical problem-solving into transparent, step-by-step workflows. By treating molecular transformations as modular "chemical operations", the framework enables slow-thinking reasoning, mirroring the logic of mathematical proofs while grounding solutions in real-world chemical constraints. We evaluate models on two high-impact tasks: Molecular Property Optimization and Chemical Reaction Prediction. These tasks mirror real-world challenges while providing structured evaluability. We further provide ChemCoTDataset, a pioneering 22,000-instance chemical reasoning dataset with expert-annotated chains of thought to facilitate LLM fine-tuning. By providing annotated trainable datasets, a reasoning taxonomy, and baseline evaluations, our work bridges the gap between abstract reasoning methods and practical chemical discovery, establishing a foundation for advancing LLMs as tools for AI-driven scientific innovation. Equal contributors, Work done as a student researcher at IDEA. Corresponding Authors, Team Leader, Connect Email: lihao1984@pku.edu.cn
@@ 1 Introduction
To address these limitations, we introduce ChemCoTBench, a step-by-step, application-oriented, and high-quality benchmark for evaluating LLM reasoning in chemical applications. A core innovation of ChemCoTBench is its formulation of complex chemical tasks, specifically targeting molecular modeling and design (Fig.1), into explicit sequences of verifiable modular chemical operations on SMILES structures (e.g., substructure addition, deletion, or substitution). This approach allows for a granular assessment of an LLM’s ability to execute and chain together fundamental chemical transformations. The benchmark features progressively challenging tasks, spanning from basic molecular understanding and editing to property-guided structure optimization and complex multi-molecule chemical reactions. High-quality evaluation is ensured through a dual validation process combining LLM judgment with expert review from 13 chemists. Furthermore, ChemCoTDataset is introduced as the first chemical reasoning dataset with precise chain-of-thought labels. Its 22,000 instances facilitate effective fine-tuning of Large Language Models.
We evaluate the chemical reasoning ability across reasoning-enhanced and non-reasoning LLMs. Experimental results reveal room for improvement in reasoning LLMs, particularly open-source and distilled-reasoning LLMs, when addressing complex chemical problems. While these models demonstrate strong performance in complex mathematical and coding tasks, they are unable to organize chemical knowledge and establish step-wise modular chemical operations due to the scarcity of chemical reasoning data. Notably, ChemCoTDataset, the large chemical CoT dataset provided by ChemCoTBench, is shown to enhance chemical reasoning performance, effectively addressing the reasoning data scarcity issue in chemical reasoning domain for LLMs.
@@ 2 Related Works
LLM Chain-of-Thoughts.  LLMs have progressed from text generators to reasoning systems, with [67]’s Chain-of-Thought enabling stepwise problem decomposition via "slow-thinking" paradigms. These reasoning-enhanced LLMs have shown impressive performance in domains requiring systematic problem-solving skills, particularly in mathematics [51], coding [27], and multi-modality tasks [69]. Models like DeepSeek-R1 [13], Gemini [59], and Anthropic Claude [56] have achieved notable results on mathematical benchmarks like MATH [19] and GSM8K [6], while also excelling at programming. Recent studies have begun exploring LLMs for chemical tasks, such as synthesis planning [4] and computational chemistry [48, 54, 26]. However, these efforts lack a systematic evaluation of LLMs’ chemical reasoning capabilities, spanning spatial reasoning, domain-specific knowledge integration, and multi-step logical inference.
@@ 3 ChemCoTBench Construction
ChemCoTBench contains 1,495 samples across 22 chemical tasks as the benchmark dataset, as shown in Fig 2(a). 22,000 high-quality samples with chain-of-thoughts annotations are further sampled to form the ChemCoTDataset. ChemCoTBench was constructed through over 1,800 hours of combined expert and LLM-assisted annotation. It comprises four main tasks and 22 subtasks, covering a broad spectrum of chemical challenges. We define the reasoning steps of each task as modular chemical operations, as shown in the bottom two lines of Fig. 3. ChemCoTBench is guided by two core principles: Diversity and Quality. Molecular diversity is ensured by systematically selecting compounds with varied scaffolds and functional groups, enabling broad coverage of real-world chemical scenarios. To ensure high data quality, all benchmark samples undergo multi-stage hybrid review by LLMs and expert chemists, with prompt templates iteratively refined to meet subtask-specific requirements.
@@ 3.3 Quality Review & Refinement
Domain Expert Models
Ether0 Failed 0.35 Failed Failed 63 94 76 78
@@ 4.2 Evaluated LLMs
Our evaluation includes three model categories: (1) Reasoning LLMs with explicit step-by-step reasoning, including Deepseek-R1 [13], o1-mini [61], o3-mini [62], Gemini-2.5-pro [58], Claude-3.7-Sonnet-thinking [56], Qwen-3-thinking [63], Llama-Nemotron-thinking [2]; (2) General-purpose non-reasoning LLMs without specialized reasoning mechanisms including GPT-4o [24], Qwen-2.5/3 [70], Llama-3.3 [12], Gemma-2 [60], Phi-4 [1], OLMo2 [47] (3) Biomolecular LLMs BioMedGPT [41], BioMistral [32], and Text+Chem T5 [5]. This comprehensive comparison evaluates whether reasoning-specific capabilities provide advantages over domain-specific models in challenging chemical reasoning tasks. Details of evaluation implementation in Appendix C.2.
@@
$\Delta$ SR% $\Delta$ SR% $\Delta$ SR% $\Delta$ SR% $\Delta$ SR% $\Delta$ SR%
@@
Gemini-2.5-pro-think -0.22 76 1.06 70 0.28 84 0.36 74 -0.02 35 0.06 68
Claude3.7-sonnet-think 0.41 80 0.37 75 0.12 73 0.17 63 0.01 49 0.02 57
DeepSeek-R1 0.47 69 0.80 80 0.17 72 0.12 62 -0.02 29 0.01 41
o3-mini@20250103 0.26 59 0.81 85 0.21 86 0.19 69 -0.03 23 0.01 45
@@
Qwen3-235B-A22B-think 0.05 40 0.20 40 0.02 24 0.03 31 -0.01 23 0.01 31
Qwen3-32B-think -0.01 1 0.13 19 0.01 9 0.0 4 -0.02 3 -0.02 6
Llama-Nemo-49B-think -0.24 7 0.25 25.2 0.10 41 0.03 29.9 -0.02 6 -0.01 11.2
@@
GPT-4o@20241120 -0.09 37 0.92 80 0.13 70 0.07 48 -0.02 30 -0.00 39
DeepSeek-V3 0.09 33 0.57 92 0.08 46 0.03 28 0.00 18 -0.01 29
Gemini-2.0-flash 0.37 72 0.28 58 0.13 79 0.15 63 -0.02 34 0.01 38
Qwen3-235B-A22B 0.03 21 0.18 45 0.07 34 0.04 26 -0.01 18 0.02 25
Qwen3-32B 0.0 2 0.08 20 0.02 14 -0.01 6 -0.02 6 -0.02 5
Qwen2.5-72B-Instruct -0.03 41 0.34 59 0.07 57 0.04 40 -0.02 26 -0.00 40
Qwen2.5-32B-Instruct 0.15 44 0.49 65 0.09 54 0.05 32 -0.02 19 0.01 31
Llama-3.1-70B-Instruct 0.02 35 0.72 81 0.15 61 -0.00 31 -0.01 30 0.01 40
Llama-Nemo-Super-49B -0.01 24 0.34 40 0.08 43 -0.00 16 -0.00 15 0.01 27
Gemma-2-27b-it 0.01 31 0.39 69 0.07 56 -0.02 15 -0.00 16 -0.00 17
Phi-4-14B -0.26 44 0.22 53 0.17 74 -0.02 18 -0.03 14 -0.00 22
OLMo2-32B-Instruct -1.71 11 1.21 46 0.08 40 -0.05 7 -0.03 8 -0.02 12
Domain Expert Models
Ether0 0.0 0 0.0 0 0.0 0 0.0 0 0.0 0 0.0 0
@@ 4.3 LLMs’ Performance on Solving ChemCoTBench
4.4 Evaluating Distillation Strategies in Chemical Reasoning
@@
Our preceding analyses underscored the critical role of advanced reasoning capabilities (or "slow thinking") for tackling complex chemical tasks. This motivates our exploration of distillation strategy [13] as a standard method to bolster this capability in open-source LLMs.
@@
Challenges in Distilling Chemical Reasoning: Distilling CoT capabilities from advanced LLMs (e.g., using DeepSeek-R1-generated samples [75, 13]) is a common strategy to enhance reasoning in smaller models. However, this approach proves significantly limited for specialized chemical reasoning. Our experiments (Fig.4) show that Qwen2.5-Instruct models distilled for CoT exhibit little to no improvement on ChemCoTBench chemical subtasks compared to their non-distilled counterparts; indeed, smaller base models (1.5B-32B) often perform comparably or better. While effective for general domains like code and math (Fig.4), this distillation strategy falters in chemistry, likely due to insufficient volume or specificity of chemical CoT samples in the distillation process, hindering the development of robust step-by-step chemical reasoning. Moreover, smaller distilled models (<7B) frequently produce lengthy, repetitive, and irrelevant (hallucinatory) thought processes. These findings suggest that direct CoT distillation, without substantial domain-specific adaptation, is an ineffective standalone method for improving chemical reasoning in open-source models. From the bottom row of Fig.4, an inverse correlation is observed between the model’s performance on Math/Code and its OOD performance in chemistry: specifically, S1.1-distill [46] outperforms R1-distill [13] on MATH500 but underperforms it on multiple chemical subtasks.
@@
Figure 4: The top two rows compare the reasoning performance of the Qwen-2.5-Instruct series against its DeepSeek-R1-distilled versions. The bottom row propose the comparison between two reasoning models with distillation strategy (R1-Distill, S1.1-Distill), and their Qwen-2.5 backbone.
Table 4: Investigation of methods to enhance chemical reasoning. We propose two approaches: prompting with chemical reasoning templates and supervised fine-tuning with ChemCoTDataset. Experimental results with Qwen-2.5 backbones (different scales from 1.5B to 32B) demonstrate that coarse guidance from reasoning templates cannot yield stable performance, while SFT with ChemCoTDataset provides significant reasoning boosting, verifying the effectiveness of ChemCoTDataset.
Models Mol-Understanding Mol-Editing
[…]
1.5B 1.32 1.17 0.07 0.15 0 0.15 0.05
1.5B-CoT-Template 0.40 1.04 0.07 0.58 0.05 0.15 0.02
1.5B-CoT-SFT 0.35 0.69 0.12 0.78 0.20 0.25 0.07
7B 0.43 1.04 0.09 0.82 0.15 0.3 0.15
7B-CoT-Template 0.25 1.21 0.09 0.57 0.15 0.45 0.15
7B-CoT-SFT 0.33 0.69 0.31 0.45 0.40 0.45 0.15
14B 0.42 1.1 0.11 0.67 0.35 0.65 0.2
14B-CoT-Template 0.35 0.91 0.12 0.62 0.3 0.4 0.2
14B-CoT-SFT 0.41 0.70 0.25 0.63 0.35 0.70 0.38
32B 0.35 0.95 0.15 0.60 0.45 0.55 0.5
32B-CoT-Template 0.33 0.74 0.12 0.70 0.4 0.65 0.4
32B-CoT-SFT 0.29 0.72 0.17 0.72 0.55 0.66 0.53
4.5 Effective Methods for Enhancing Chemical Reasoning
Given the limitations of direct distillation, we explore effective strategies to enhance chemical reasoning capabilities. We investigate two approaches: prompting with chemical reasoning templates and supervised fine-tuning (SFT) with our ChemCoTDataset.
Prompt engineering cannot bring stable chemical reasoning improvements: we first evaluate the CoT prompting strategy: providing only coarse strategic guidance (CoT templates). The results in Table. 4 demonstrate that the prompting strategy cannot yield stable and significant performance gains across all chemical reasoning tasks. Specifically, for the functional-group detection task, models from 1.5B to 32B show stable improvements. However, for other chemical reasoning tasks, CoT prompting strategy shows unstable influence due to the lack of chemical knowledge.
SFT on ChemCoTDataset boosts chemical reasoning: We further explored enhancing chemical reasoning using our high-quality, domain-specific ChemCoTDataset via supervised funetuning. This dataset was meticulously curated to minimize hallucinations and align with expert thought processes, which we posited would be vital for chemical reasoning tasks. We test this by evaluating the SFT strategy augmented with detailed step-by-step reasoning processes from our dataset. The results in Table. 4 consistently demonstrate that our large-scale chemical CoT dataset significantly enhances the chemical reasoning capabilities of Qwen-2.5 models across various scales (1.5B to 32B) when used in this way. Augmentation with SFT processes yielded stable and substantial performance gains across all evaluated tasks.